# Spiking Meets Attention: Efficient Remote Sensing Image Super-Resolution with Attention Spiking Neural Networks

### v2

@@ III-B Deformable Similarity Attention
where we adopt the Gumbel-Softmax [54] to achieve the non-differentiable $\rm argmax$ function, and $E$ means the operation of average pooling and feature reshaping.
@@
Finally, the fused features are summarized with the original features $\mathbf{F}$ and fed into the multilayer perceptron (MLP), which contains two fully-connected layers, to obtain the final output:
@@
TABLE I: Quantitative comparison of SpikeSR with state-of-the-art methods on three remote sensing datasets. FLOPs are measured corresponding to an LR image of $160\times 160$ pixels.
Methods #Param. FLOPs AID [55] DOTA [56] DIOR [57] Average
@@
SRCNN [19] 20K 0.512G 29.70 0.7741 32.10 0.8264 29.49 0.7768 30.43 0.7924
VDSR [20] 667K 17.08G 30.44 0.8004 33.22 0.8569 30.36 0.8036 31.34 0.8203
EDSR [21] 1518K 50.77G 30.65 0.8086 33.64 0.8648 30.63 0.8116 31.64 0.8283
CARN [28] 1112K 40.39G 30.66 0.8068 33.66 0.8633 30.64 0.8102 31.65 0.8268
IMDN [29] 715K 18.18G 30.71 0.8076 33.70 0.8641 30.73 0.8115 31.71 0.8277
RFDN-L [58] 681K 16.49G 30.69 0.8074 33.73 0.8642 30.72 0.8114 31.71 0.8277
LatticeNet [59]  [^7] K 19.39G 30.73 0.8089 33.75 0.8653 30.75 0.8126 31.74 0.8289
HNCT [60] 364K 8.48G 30.79 0.8104 33.83 0.8664 30.80 0.8136 31.81 0.8301
FMEN [30] 1046K 26.72G 30.65 0.8063 33.66 0.8631 30.66 0.8104 31.66 0.8266
RLFN [61] 544K 13.25G 30.70 0.8074 33.69 0.8636 30.70 0.8110 31.70 0.8273
ESRT [62] 752K 26.06G 30.77 0.8102 33.75 0.8668 30.81 0.8142 31.78 0.8304
SwinIR-light [27] 897K 23.56G 30.83 0.8114 33.94 0.8677 30.85 0.8149 31.87 0.8313
Omni-SR [63] 2803K 70.98G 30.89 0.8142 33.94 0.8695 30.89 0.8170 31.91 0.8336
NGswin [64] 995K 12.73G 30.79 0.8107 33.87 0.8667 30.79 0.8140 31.82 0.8305
SPIN [31]  [^5] K 18.91G 30.78 0.8098 33.85 0.8673 30.82 0.8139 31.82 0.8303
HiT-SR [32] 792K 21.04G 30.87 0.8138 33.93 0.8689 30.89 0.8167 31.90 0.8331
@@ III-C Fusion Block
where $\rm TA$ and $\rm SA$ denote temporal and spatial attention [44], $\sigma$ means a sigmoid function and $\otimes$ denotes feature multiplication. The final output of FB is obtained by summing $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$.
TABLE II: Quantitative comparison of SpikeSR with state-of-the-art methods on 30 scene types of AID datasets.
Scene types EDSR [21] CARN [28] IMDN [29] ESRT [62] SwinIR [27] SPIN [31] HiT-SR [32] SpikeSR
@@
Figure 5: Qualitative comparison of state-of-the-art efficient models for ×4 SR task on AID test set.
Figure 6: Qualitative comparison of state-of-the-art efficient SR models for ×4 SR task on DOTA dataset.
Figure 7: Qualitative comparison of state-of-the-art efficient SR models for ×4 SR task on DIOR dataset.
IV Experiments
Datasets. We use the AID dataset [55] as the training set, a large-scale remote sensing benchmark for scene classification, consisting of 30 different scene categories. The AID dataset includes 10,000 HR images, where we randomly select 3,000 for training and 900 for validation. The LR samples are generated by bicubic downsampling. Following TTST [65], we also evaluate our method on the DOTA [56] and DIOR [57] datasets, which contain 900 and 1,000 images, respectively.
Implementation Details. During model training, the learning rate is fixed to $10^{-4}$, and the training procedure stops after 1000 epochs with a batch size of 4. Adam optimizer is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Data augmentation includes random rotations of 90°, 180°, 270°, and horizontal flips on 64×64 patches. The channel number, embedding dimension of cross-attention, and MLP rate of small, medium, and final SpikeSR are set to $\{40,24,72\}$, $\{56,24,72\}$, and $\{64,32,100\}$, respectively. The number of SAGs is 4, with 2 SABs in each SAG. Time step is set to $\rm T=4$.
Metrics. We use the widely used peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) to evaluate SR performance. The results are measured on the Y channel after converting RGB to YCbCr space. For a fair comparison, comparative methods are retained from scratch on the AID dataset, adhering to their official implementation settings.
IV-A Comparison with Efficient Models
We compare our SpikeSR with state-of-the-art (SOTA) efficient SR methods, including CNN-based models of CARN [28], LatticeNet [59], RLFN [61], etc, and transformer-based approaches of SwinIR [27], SPIN [31], HiT-SR [32], etc. We also report results for the small and medium versions of our SpikeSR, denoted as SpikeSR-S and Spike-M, respectively.
Quantitative comparisons. The quantitative results of various methods are reported in Table I. We can observe that our SpikeSR achieves the best performance across three benchmarks, outperforming SOTA CNN- and transformer-based SR models. For example, on AID, DOTA, and DIOR datasets, SpikeSR improves PSNR by 0.08 dB, 0.04 dB, and 0.1 dB, respectively, compared to the impressive SwinIR. Moreover, the small version of SpikeSR requires fewer parameters (472K vs. 897K) and FLOPs (15.21G vs. 23.56G) than SwinIR, yet achieves superior average performance.
@@ IV-A Comparison with Efficient Models
V Ablation Study
We conduct ablation studies to assess the impact of key components in SpikeSR. The experimental results in Table III are measured on the AID-tiny dataset [65]. In particular, the Baseline model is constructed by removing the HDA and DSA and replacing them with standard temporal attention (TA), channel attention (CA), and spatial attention (SA) mechanisms. For a fair comparison, we increase the number of $m$, $n$, and channel dimensions to 8, 4, and 256, respectively, which ensures a similar number of parameters with our SpikeSR. Similarly, those settings of Variant-A are modified to 10, 5, and 128, respectively.
TABLE III: Ablation on different variants of our SpikeSR.
Methods TA CA SA HAD DSA #Param. PSNR
@@ V Ablation Study
TABLE IV: Ablation on feature pyramid and deformable convolution.
@@
Figure 8: Feature visualizations. The feature obtained by HDA is sharper and preserves more details, indicating high-quality feature representation.
TABLE V: Performance and complexity analysis of SpikeSR with different numbers of blocks.
Number of $m$ 2 3 4 5 6
@@
Figure 9: Analysis of local attribution maps (LAMs) [66] and diffusion index (DI). The proposed DAS helps SpikeSR exploit more useful information against CARN and HiT-SR.
Effectiveness of HDA and DSA. Table III indicates how the SR performance is influenced by the HDA and DSA. Comparing the PSNR values of the Baseline and Variant-A reveals that HDA contributes a 0.12 dB improvement. This suggests that enhancing the correlation between the temporal and channel dimensions delivers better recalibration of the membrane potentials. More intuitively, we visualize the feature maps to highlight the impact of HDA, as shown in Fig. 8. The results illustrate that HDA effectively refines the feature representation by emphasizing salient details and suppressing irrelevant information.
By introducing the proposed DSA to the Baseline model, the PSNR results can be improved by a large margin of 0.21 dB. When employing HDA and DSA simultaneously, the resulting SpikeSR achieves an additional 0.08 dB improvement compared to Variant-B. To better demonstrate its effectiveness in capturing global self-similarity priors, we provide the LAM visualization and diffusion index in Fig. 9. As observed, DSA generates more pronounced LAMs and significantly increases the DI, indicating that the model activates more valuable pixels for accurate SR.
Feature Pyramid and DConv. Due to the scale diversity of objects in RSIs, we constructed a feature pyramid to grasp the self-similarity in multiple levels. As listed in Table IV, the use of feature pyramid improves the performance by 0.05 dB without introducing additional parameters. Furthermore, to demonstrate the effectiveness of deformable convolution, we remove this component, which leads to a severe performance drop by 0.12 dB. This illustrates that the self-similar patches contain massive irrelevant and misaligned contents, and direct fusion may introduce interference, thus resulting in suboptimal performance.
Network Depth. We evaluate the impact of the network depth by changing the number of SAGs of our SpikeSR from 2 to 6 blocks. As reported in Table V, SpikeSR achieves the highest SR performance when $m=3$. While increasing the number of $m$ may further improve the reconstruction, it also brings larger model size. Therefore, we set $m=4$ in our final model, considering the trade-off between performance and computational complexity.
VI Conclusion
In this paper, we investigate the application of SNNs for efficient SR of remote sensing images. Motivated by the observation that LIF neurons exhibit a higher spike rate in degraded images, we integrate SNNs with convolutions for improved feature representation. Besides, a hybrid dimension attention is employed to modulate the spike response, further refining salient information. To incorporate valuable prior knowledge for more accurate SR, we propose a deformable similarity attention module, capturing global self-similarity across multiple feature levels. Extensive experiments on various remote sensing datasets demonstrate the efficacy and effectiveness of the proposed SpikeSR model.

### v3

@@ 2.2 Deformable Similarity Attention
where $\mathbf{\bar{P}}_{i}$} is the patch in $\mathbf{\bar{F}}$}, and $E$ means the operation of average pooling and feature reshaping. We adopt the Gumbel-Softmax wang2021exploring to achieve the non-differentiable $\rm argmax$ function.
@@
Finally, the fused features are summarized with the original features $\mathbf{F}$ and fed into the multilayer perceptron (MLP) to obtain the final output:
@@
Table 1: Quantitative comparison of SpikeSR with state-of-the-art methods on three remote sensing datasets. FLOPs are measured corresponding to an LR image of $160\times 160$ pixels. Note that we set $\rm{T}=1$ to evaluate the model complexity of SpikeSR for fair comparison.
Methods #Param. FLOPs AID aid DOTA dota DIOR dior Average
@@
SRCNN srcnn 20K 0.512G 29.70 0.7741 32.10 0.8264 29.49 0.7768 30.43 0.7924
VDSR vdsr 667K 17.08G 30.44 0.8004 33.22 0.8569 30.36 0.8036 31.34 0.8203
EDSR edsr 1518K 50.77G 30.65 0.8086 33.64 0.8648 30.63 0.8116 31.64 0.8283
CARN carn 1112K 40.39G 30.66 0.8068 33.66 0.8633 30.64 0.8102 31.65 0.8268
IMDN imdn 715K 18.18G 30.71 0.8076 33.70 0.8641 30.73 0.8115 31.71 0.8277
RFDN-L rfdn 681K 16.49G 30.69 0.8074 33.73 0.8642 30.72 0.8114 31.71 0.8277
LatticeNet latticenet  [^7] K 19.39G 30.73 0.8089 33.75 0.8653 30.75 0.8126 31.74 0.8289
HNCT hnct 364K 8.48G 30.79 0.8104 33.83 0.8664 30.80 0.8136 31.81 0.8301
FMEN fmen 1046K 26.72G 30.65 0.8063 33.66 0.8631 30.66 0.8104 31.66 0.8266
RLFN rlfn 544K 13.25G 30.70 0.8074 33.69 0.8636 30.70 0.8110 31.70 0.8273
ESRT esrt 752K 26.06G 30.77 0.8102 33.75 0.8668 30.81 0.8142 31.78 0.8304
SwinIR-light swinir 897K 23.56G 30.83 0.8114 33.94 0.8677 30.85 0.8149 31.87 0.8313
Omni-SR omnisr 2803K 70.98G 30.89 0.8142 33.94 0.8695 30.89 0.8170 31.91 0.8336
NGswin ngswin 995K 12.73G 30.79 0.8107 33.87 0.8667 30.79 0.8140 31.82 0.8305
SPIN spin  [^5] K 18.91G 30.78 0.8098 33.85 0.8673 30.82 0.8139 31.82 0.8303
HiT-SR hitsr 792K 21.04G 30.87 0.8138 33.93 0.8689 30.89 0.8167 31.90 0.8331
@@ 2.3 Fusion Block
where $\rm TA$ and $\rm SA$ denote temporal and spatial attention yao2023attention , $\sigma$ means a sigmoid function and $\otimes$ denotes feature multiplication. The final output of FB is obtained by summing $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$.
Table 2: Quantitative comparison of SpikeSR with state-of-the-art methods on 30 scene types of AID datasets.
Scene types EDSR edsr CARN carn IMDN imdn ESRT esrt SwinIR-L swinir SPIN spin HiT-SR hitsr SpikeSR
@@
Figure 4: Qualitative comparison of SOTA efficient models for $\times$4 SR task on AID test set.
Figure 5: Qualitative comparison of SOTA efficient SR models for $\times$4 SR task on DOTA dataset.
Figure 6: Qualitative comparison of SOTA efficient SR models for $\times$4 SR task on DIOR dataset.
3 Experiments
Datasets. We use the AID dataset aid as the training set, a large-scale remote sensing benchmark for scene classification, consisting of 30 different scene categories. The AID dataset includes 10,000 HR images, where we randomly select 3,000 for training and 900 for validation. The LR samples are generated by bicubic downsampling. Following TTST ttst , we also evaluate our method on the DOTA dota and DIOR dior datasets, which contain 900 and 1,000 images, respectively.
Implementation Details. During model training, the learning rate is fixed to $10^{-4}$, and the training procedure stops after 1000 epochs with a batch size of 4. Adam optimizer is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Data augmentation includes random rotations of 90°, 180°, 270°, and horizontal flips on $64\times 64$ patches. The channel number, embedding dimension of cross-attention, and MLP rate of small, medium, and final SpikeSR are set to $\{40,24,72\}$, $\{56,24,72\}$, and $\{64,32,100\}$, respectively. The number of SAGs is 4, with 2 SABs in each SAG. Time step is set to $\rm T=4$.
Metrics. The widely used peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) are used to evaluate SR performance. The results are measured on the Y channel after converting RGB to YCbCr space. For a fair comparison, all SR methods are retained on an RTX 4090 GPU from scratch using the AID dataset, adhering to their official implementation settings.
3.1 Comparison with Efficient Models
We compare our SpikeSR with state-of-the-art (SOTA) efficient SR methods, including CNN-based models of CARN carn , LatticeNet latticenet , RLFN rlfn , etc, and transformer-based approaches of SwinIR swinir , SPIN spin , HiT-SR hitsr , etc. We also report results for the small and medium versions of our SpikeSR, denoted as SpikeSR-S and Spike-M, respectively.
Quantitative comparisons. The quantitative results of various methods are reported in Table 1. We can observe that our SpikeSR achieves the best performance across three benchmarks, outperforming SOTA CNN- and transformer-based SR models. For example, on AID, DOTA, and DIOR datasets, SpikeSR improves PSNR by 0.08 dB, 0.04 dB, and 0.1 dB, respectively, compared to the impressive SwinIR. Moreover, the small version of SpikeSR requires fewer parameters (472K vs. 897K) and FLOPs (15.21G vs. 23.56G) than SwinIR, yet achieves superior average performance.
@@ 3.1 Comparison with Efficient Models
4 Ablation Study
We conduct ablation studies to assess the impact of key components in SpikeSR. The experimental results in Table 3 are measured on the AID-tiny dataset ttst . In particular, the Baseline model is constructed by removing the HDA and DSA and replacing them with standard temporal attention (TA), channel attention (CA), and spatial attention (SA) mechanisms. For a fair comparison, we increase the number of $m$, $n$, and channel dimensions to 8, 4, and 256, respectively, which ensures a similar number of parameters with our SpikeSR. Similarly, those settings of Variant-A are modified to 10, 5, and 128, respectively.
Figure 7: (a) Analysis of local attribution maps (LAMs) lam and diffusion index (DI). The proposed DAS helps SpikeSR exploit more useful information against CARN and HiT-SR. (b) Feature visualizations. The feature obtained by HDA is sharper and preserves more details, indicating high-quality feature representations.
Effectiveness of HDA and DSA. Table 3 indicates how the SR performance is influenced by the HDA and DSA. Comparing the PSNR values of the Baseline and Variant-A reveals that HDA contributes a 0.12 dB improvement. This suggests that enhancing the correlation between the temporal and channel dimensions delivers better recalibration of the membrane potentials. More intuitively, we visualize the feature maps to highlight the impact of HDA, as shown in Fig. 7(b). The results illustrate that HDA effectively refines the feature representation by emphasizing salient details and suppressing irrelevant information.
By introducing the proposed DSA to the Baseline model, the PSNR results can be improved by a large margin of 0.21 dB. When employing HDA and DSA simultaneously, the resulting SpikeSR achieves an additional 0.08 dB improvement compared to Variant-B. To better demonstrate its effectiveness in capturing global self-similarity priors, we provide the LAM visualization and diffusion index in Fig. 7(a). As observed, DSA generates more pronounced LAMs and significantly increases the DI, indicating that the model activates more valuable pixels for accurate SR.
Table 3: Ablation on different variants of our SpikeSR.
Methods TA CA SA HDA DSA #Param. PSNR (dB)
@@ 4 Ablation Study
Table 4: Ablation on feature pyramid and deformable convolution.
@@
Table 5: Performance and complexity analysis of SpikeSR with different numbers of blocks.
Blocks 2 3 4 5 6
@@
Feature Pyramid and DConv. Due to the scale diversity of objects in RSIs, we constructed a feature pyramid to grasp the self-similarity in multiple levels. As listed in Table 4, the use of feature pyramid improves the performance by 0.05 dB without introducing additional parameters. Furthermore, to demonstrate the effectiveness of deformable convolution, we remove this component, which leads to a severe performance drop by 0.12 dB. This illustrates that the self-similar patches contain massive irrelevant and misaligned contents, and direct fusion may introduce interference, thus resulting in suboptimal performance.
Network Depth. We evaluate the impact of the network depth by changing the number of SAGs of our SpikeSR from 2 to 6 blocks. As reported in Table 4, SpikeSR achieves the highest SR performance when $m=3$. While increasing the number of $m$ may further improve the reconstruction, it also brings larger model size. Therefore, we set $m=4$ in our final model, considering the trade-off between performance and computational complexity.
5 Conclusion and Limitation
In this paper, we investigate the application of SNNs for efficient SR of remote sensing images. Motivated by the observation that LIF neurons exhibit a higher spike rate in degraded images, we integrate SNNs with convolutions for improved feature representation. Besides, a hybrid dimension attention is employed to modulate the spike response, further refining salient information. To incorporate valuable prior knowledge for more accurate SR, we propose a deformable similarity attention module, capturing global self-similarity across multiple feature levels. Extensive experiments on various remote sensing datasets demonstrate the efficacy and effectiveness of the proposed SpikeSR model. We believe our exploration can facilitate the practical application of energy-efficient models in remote sensing area.
SpikeSR still has some limitations. First, its representational capacity remains improvable. Second, the reliance on attention mechanisms introduces additional computational overhead, limiting efficiency. We leave these issues for future exploration.
Acknowledgment. This work is supported by the National Natural Science Foundation of China under Grant 423B2104 and 623B2080.